# Boundary layers, transport and universal distribution in boundary driven active systems

### v2

@@ 4.1.1 Spectrum in the large-$L$ limit
In the limit of large system size, we expect that the eigenvalues are very close to zero and therefore from Eq. (24), we get,
@@
The finite value of $k_{a}$ renders $e^{-k_{a}\,L}\approx 0$ in Eqs. (41)-(42). Further, from Eq. (27) one can find that, $a^{(3)}\approx 1,~{}a^{(1)}\approx-\frac{1+v/\sqrt{2\omega D_{e}}}{1-v/\sqrt{2\omega D_{e}}}$, and the ratio in the rhs of the Eqs. (41)-(42) takes the value, $\frac{a^{(3)}-a^{(1)}}{a^{(3)}a^{(1)}-1}\approx-1$. Consequently, we have, for $o_{x}=\pm 1$,
@@
To evaluate the subleading behaviour we need to keep the first correction terms in $\lambda$ in Eqs. (41)-(42). Since $k_{a}$ is finite, in the large $L$ limit we shall neglect $e^{-k_{a}L}$ and therefore the l.h.s. in both the equations is $e^{-k_{b}L}\approx e^{-L\sqrt{\lambda/D_{e}}}$. On the other hand, for small $\lambda$, $a^{(3)}\approx 1-\frac{v}{\omega\sqrt{D_{e}}}\sqrt{\lambda}$, whereas $a^{(1)}\approx-\frac{1+v/\sqrt{2\omega D_{e}}}{1-v/\sqrt{2\omega D_{e}}}+O(\lambda)$. This implies that the corrections in the rhs of Eqs. (41)-(42) are of $O(\sqrt{\lambda})$, and the equations to the first subleading order reads,
@@
One can solve the above equation perturbatively by considering $\lambda_{n}=\lambda_{n}^{(0)}+\lambda_{n}^{(1)}$, where $\lambda_{n}^{(0)}=-\frac{n^{2}\pi^{2}D_{e}}{L^{2}}$ are the solution at the leading order in $L$ and $\lambda_{n}^{(1)}\ll\lambda_{n}^{(0)}$ are the corrections. We want to find the leading $L$-dependence of $\lambda_{n}^{(1)}$. Putting $\lambda_{n}=\lambda_{n}^{(0)}+\lambda_{n}^{(1)}$ in Eq. (47), we get,
$e^{-L\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{D_{e}}}}\approx\mp\left(1+\frac{v^{2}}{\omega D_{e}}\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{2\omega}}\right).$ (48)
To the first subleading order the left hand side of the Eq. (48) can be written as,
$e^{-k_{b}L}\approx e^{-L\sqrt{\frac{\lambda_{n}^{(0)}+\lambda_{n}^{(1)}}{D_{e}}}}$ $=$ $e^{-L\sqrt{\frac{\lambda_{n}^{(0)}}{D_{e}}}\,\left(1+\frac{\lambda_{n}^{(1)}}{\lambda_{n}^{(0)}}\right)^{1/2}}\approx e^{-L\sqrt{\frac{\lambda_{n}^{(0)}}{D_{e}}}}~{}e^{-L\sqrt{\frac{\lambda_{n}^{(0)}}{D_{e}}}\,\left(\frac{\lambda_{n}^{(1)}}{2\,\lambda_{n}^{(0)}}\right)}$ (49)
Now, using in the above the expression for $\lambda_{n}^{(0)}$ from Eq. (45), the l.h.s. of Eq. (48) is obtained as,
$e^{-k_{b}L}\approx\mp 1\times\exp\left(-i\,\frac{n\pi}{2}\,\frac{\lambda_{n}^{(1)}}{\lambda_{n}^{(0)}}\right)\approx\mp 1\pm i\,\frac{n\pi}{2}\,\frac{\lambda_{n}^{(1)}}{\lambda_{n}^{(0)}},$ (50)
Similarly, we find the right hand side of Eq. (48) to the first subleading order,
Then, equating the l.h.s. and r.h.s. of Eq. (48), we get,
$\lambda_{n}^{(1)}\approx-\frac{2\,v^{2}}{\omega\,\sqrt{2\omega D_{e}}}\frac{\lambda_{n}^{(0)}}{L},$ (51)
and consequently, the band near $\lambda=0$ takes the form,
@@
We have already mentioned that $-2\omega$ is an eigenvalue of the evolution operator that belongs to the symmetric sector. To calculate the band of eigenvalues around it, let us consider $\lambda=-2\omega+\varepsilon$, where $\varepsilon$ is presumably small in the large $L$ limit. In this limit,
$k_{a}\approx\sqrt{\frac{v^{2}}{D^{2}}+\frac{\lambda}{D}},~{}k_{b}\approx\sqrt{\varepsilon/D}$
@@
Since $\varepsilon\to 0$, $a^{(3)}\approx-1$. In the large $L$-limit, $e^{-k_{a}L}\to 0$. Therefore, Eqs (42)-(43) reduce to
@@ 4.2 Crossover behaviour of the relaxation time
Figure 6: Crossover behaviour of the relaxation rate for different system sizes. Here, $D=1.0,v=1.0,\omega=0.1$. Points are obtained by numerically solving Eq. (41) (symmetric sector) and identifying the solution closest to zero using Mathematica. The crossover to $L$-dependent relaxation is seen to occur for system sizes very close to $L_{p}=\frac{v}{\omega}=10.0$. The large $L$ behaviour is consistent with Eq. (52) for $n=1$.
For passive particles the relaxation time because of diffusion is proportional to $L^{2}/D$. RTPs on the other hand also carries a tumble dynamics which relaxes in a timescale $\tau_{t}\sim(2\omega)^{-1}$. When the system size is much larger than the persistent length $L_{c}\sim v/\omega$, several tumble events occur before reaching the steady state and the relaxation dynamics becomes effectively passive-like. In this regime the bands are also well-separated. As $L$ is reduced, the corrections due to activity (Eq. (52)) become significant and a deviation from diffusive behaviour shows up. When the $L\sim L_{c}$, the two bands would mix and for $L<L_{c}$ they cross, implying that the diffusive dynamics now relaxes faster compared to the tumble dynamics and $\lambda=-2\omega$ determines the relaxation of the system. This behaviour is shown in Figure 6. We expect this to be a generic feature of finite-sized active systems, and more generally of processes with coloured noise.
@@ 4.3 Eigenfunctions and the large-time distribution
Once we have the spectrum, we can completely solve for the coefficients $\alpha_{n}^{i},i=(1,\dots,4)$. For $o_{x}=\pm 1$, the coefficients are related as follows:
$\alpha^{(1)}$ $=$ $-\,\alpha^{(3)}\,\frac{a^{(3)}\pm e^{k_{b}\,L}}{a^{(1)}\pm e^{k_{a}\,L}},$ (55)
$\alpha^{(2)}$ $=$ ±$\pm\,\alpha^{(1)}a^{(1)}\,e^{k_{a}\,L},$ (56)
$\alpha^{(4)}$ $=$ ±$\pm\,\alpha^{(3)}a^{(3)}\,e^{k_{b}\,L},$ (57)
and $\alpha^{(3)}$ is left arbitrary at this point. After incorporating the coefficients in Eq. (28), the expression of the eigenfunctions $|\phi_{n}(x)\rangle$ becomes,
$|\phi_{n}(x)\rangle=\alpha_{n}^{(3)}\bigg{[}\frac{a^{(3)}\pm e^{k_{b}\,L}}{a^{(1)}\pm e^{k_{a}\,L}}\left(e^{k_{a}\,x}\begin{bmatrix}a^{(1)}\\ 1\end{bmatrix}\pm e^{k_{a}(L-x)}\begin{bmatrix}1\\ a^{(1)}\end{bmatrix}\right)-\left(e^{k_{b}\,x}\begin{bmatrix}a^{(3)}\\ 1\end{bmatrix}\pm e^{k_{b}(L-x)}\begin{bmatrix}1\\ a^{(3)}\end{bmatrix}\right)\bigg{]}.$ (58)
The set of equations (55)-(57) can be simplified and the coefficients can be determined order by order in the large $L$ limit. We shall present results for the band near $\lambda=0$ and keep the terms which are $\sim$ $O(L^{-1})$.
To this order we find, for $o_{x}=\pm 1$, $e^{k_{b}\,L}\approx\mp(1-\frac{n\,\pi}{L}\,\frac{v^{2}}{\omega\,\sqrt{2\omega D_{e}}})$ using Eqs. (50)-(51), and $\alpha^{(1)}=-\,\alpha^{(3)}\,\frac{a^{(3)}\pm\,e^{k_{b}\,L}}{a^{(1)}\pm\,e^{k_{a}\,L}}\approx\mp\,\alpha^{(3)}\,[a^{(3)}\pm e^{k_{b}\,L}]\,e^{-k_{a}\,L}$ from Eq. (55). In Eq. (27), keeping the first subleading term in the expression for $a^{(3)}$ and using Eq. (52) to determine $k_{b}=\sqrt{\lambda/D_{e}}$ to $O(L^{-1})$, we find, $a^{(3)}\approx 1-vk_{b}/\omega\approx 1-i\,\frac{v}{\omega}\,\frac{n\pi}{L}$. Putting these together in Eq. (55) we obtain,
$\alpha^{(1)}=\pm\,i\,\alpha^{(3)}\,\frac{n\pi}{L}\,\frac{v}{\omega}\big{(}1-\frac{v}{\sqrt{2\omega D_{e}}}\big{)}\,e^{-k_{a}\,L}\,,{~{}\rm for~{}}n={\rm odd/even}.$ (59)
Using the above in Eq. (56) we find,
$\alpha^{(2)}=\,i\,\alpha^{(3)}\,\frac{n\pi}{L}\,a^{(1)}\,\frac{v}{\omega}\big{(}1-\frac{v}{\sqrt{2\omega D_{e}}}\big{)};$ (60)
and finally, from Eq. (57)
$\alpha^{(4)}=-\alpha^{(3)}a^{(3)}\big{(}1-i\,\frac{n\pi}{L}\,\frac{v^{2}}{\omega\sqrt{2\omega D_{e}}}\big{)}.$ (61)
Using the coefficients in Eq. (29), we find, at large times $t\gg\,\omega^{-1}$,
$|P_{\rm tr}(x,t)\rangle$$\approx$ $\sum_{n=1}^{\infty}c_{n}\,e^{\lambda_{n}t}\alpha_{n}^{(3)}\bigg{(}2\,i\,\sin\frac{n\pi x}{L}\begin{bmatrix}1\\ 1\end{bmatrix}-\,i\,\frac{n\pi v}{\omega\,L}\cos\frac{n\pi x}{L}\begin{bmatrix}1\\ -1\end{bmatrix}$ (62)
@@
$+\,i\,\frac{n\pi v}{\omega\,L}\,(1-\frac{v}{\sqrt{2\omega\,D_{e}}})\,\bigg{\{}(-1)^{n-1}e^{-k_{a}(L-x)}\begin{bmatrix}a_{n}^{(1)}\\ 1\end{bmatrix}+e^{-k_{a}\,x}\begin{bmatrix}1\\ a_{n}^{(1)}\end{bmatrix}\bigg{\}}\bigg{)},~{}~{}~{}~{}~{}~{}~{}~{}~{}$
with $\lambda_{n}\approx-\frac{n^{2}\pi^{2}D_{e}}{L^{2}},~{}n>0$; here we have used $a_{n}^{(1)}\,a_{n}^{(2)}=1$. Each term of the summation in Eq. (62) is an eigenfunction evaluated to $O(L^{-1})$. In the following we shall use $\alpha_{n}$ instead of $\alpha_{n}^{(3)}$. To the leading order, each of the eigenfunctions is simply proportional to $\sin\frac{n\pi x}{L}$ and the probability density takes a passive-like form with an effective diffusion constant $D_{e}$:
$\lim_{t\rightarrow\infty}\lim_{L\rightarrow\infty}P(x,t)\approx\sum_{n=1}^{\infty}c_{n}\alpha_{n}\,e^{\lambda_{n}t}4\,i\,\sin\frac{n\pi x}{L}\equiv P_{\rm passive}(x,t)|_{D\rightarrow D_{e}},$ (63)
while the magnetisation vanishes. In particular, the relaxation mode is given by $\phi_{1}(x)\propto\sin\frac{\pi x}{L}$. Note that the active contributions in excess to the effective passivelike contributions, proportional to the active speed $v$, appear at $O(L^{-1})$ in each term of the distribution $P_{\rm tr}(x,t)=[1~{}~{}1]\,|P_{\rm tr}(x,t)\rangle$ and magnetization $Q_{\rm tr}(x,t)=[1~{}-1]\,|P_{\rm tr}(x,t)\rangle$. Thus at very large times when only very few $n$’s contribute, the active part in the distribution and the magnetisation appears only at the subleading order. However, this is not true in the large but intermediate times when larger $n$’s also contribute.
@@
$+~{}m_{0}\,\frac{n\pi v}{L\omega}\,\cos\frac{n\pi x_{0}}{L}\sin\frac{n\pi x}{L}\begin{bmatrix}1\\ 1\end{bmatrix}-\frac{n\pi v}{L\omega}\,\sin\frac{n\pi x_{0}}{L}\cos\frac{n\pi x}{L}\begin{bmatrix}1\\ -1\end{bmatrix}\bigg{)}.$
At relaxation time scales ($t\gtrsim L^{2}/D_{e}$), only the small $n$ terms ($n=1,2$ etc.) survive. Here the dominating contribution comes from the first term of Eq. (64) which is just the effective passive expression, while the active contribution occurs as subleading corrections. However, at intermediate times $\omega^{-1}\ll t\ll L^{2}/D_{e}\,$, all the $n$’s contribute and it is not obvious whether the active contributions constitute only a correction to the effective passive behaviour. For this we aim to evaluate the summations in a closed form. Generally this is not doable for the summations at hand, but we can make progress for large values of $L$ using the Poisson summation formula and the symmetries of the distribution. The details and a closed form expression of $|P_{\rm tr}(x,t)\rangle$ is given in Appendix B. For finite $x_{0}\ll L$ the system is well approximated by a semi-infinite line, for which we discuss the results below.
@@ 4.4 Semi-infinite line with absorbing barrier at $x=0$:
On the semi-infinite geometry the summations in the Eq. (64) can be converted to integrals and the state vector at large times take the form:
@@
$+\frac{v}{\omega}\,\frac{\pi}{(4\pi D_{e}t)^{3/2}}\left\{(x-x_{0})\,e^{-\frac{(x-x_{0})^{2}}{4D_{e}t}}-(x+x_{0})\,e^{-\frac{(x+x_{0})^{2}}{4D_{e}t}}\right\}\begin{bmatrix}1\\ -1\end{bmatrix}\bigg{)}.$
The same expression is obtained by taking the limit $L\gg x,x_{0}$ in Eq. (B.7). All the terms in the above expression are $\sim O(t^{-3/2})$, and therefore ‘passive’ as well as the ‘active’ contributions are equally important in determining the large time behaviour in the semi infinite domain. From Eq. (65) we find the distribution and magnetisation profile by respectively adding and subtracting the rows:
@@
The profiles carry an interesting structure. In each of Eqs. (66)-(67), the term that falls exponentially in distance represents the kinetic boundary layer with a ‘skin depth’ $k_{a}^{-1}$, identical to that in the steady state. The other terms form the ‘scaling’ part of the profile. In Eq. (66) the last but one term falls rapidly if the initial position exceeds the skin depth.
@@
which is finite throughout the system for any finite $x_{0}$, and can actually attain dominant role for highly persistent (small $\omega$ or large $v$) run and tumble motion. The effect of activity becomes more prominent when the particle is in the ‘boundary layer’ $x\lesssim k_{a}^{-1}$. Beyond that this part decays slowly as $x^{-1}$, and asymptotically the contribution from initial position $x_{0}$ only remains. In fact, it can be shown from Eq. (66) that, when both $x$ and $t$ is very large such that $u=x/\sqrt{D_{e}t}$ is finite, both the active and passive contributions are $\sim(D_{e}t)^{-1}$, and their ratio becomes,
$R(x_{0})=\lim_{\{x,t\rightarrow\infty,\frac{x}{\sqrt{D_{e}t}}=u\}}\frac{P^{\infty}_{\rm active}}{P^{\infty}_{\rm passive}}=\frac{v}{2\,\omega}\big{(}m_{0}+\frac{v}{\sqrt{2\,\omega\,D_{e}}}\big{)}\frac{1-e^{-k_{a}x_{0}}}{x_{0}}.$ (69)
Figure 8: $R(x_{0})$ from Eq. (69) as a function of $1/\omega$ for (a) $m_{0}=-1$, (b) $m_{0}=0$, (c) $m_{0}=1$ for different values of $x_{0}$. Here $v=D=1$.

### v4

@@ 4.2.1 Spectrum in the large-$L$ limit
In the limit of large system size, we expect that the distribution takes very long to relax to the steady state and the particle dynamics is close to diffusion. Consequently the low lying $\lambda$ values are close to zero and from Eq. (24), we get,
@@
The finite value of $k_{a}$ renders $e^{-k_{a}\,L}\approx 0$ in Eqs. (41)-(42). Further, from Eq. (27) one can find that, $a^{(3)}\approx 1,~a^{(1)}\approx-\frac{1+v/\sqrt{2\omega D_{e}}}{1-v/\sqrt{2\omega D_{e}}}$, and the ratio in the r.h.s. of the Eqs. (41)-(42) takes the value, $\frac{a^{(3)}-a^{(1)}}{a^{(3)}a^{(1)}-1}\approx-1$. Consequently, we have, for $o_{x}=\pm 1$,
@@
To evaluate the subleading behaviour we need to keep the first correction terms in $\lambda$ in Eqs. (41)-(42). Since $k_{a}$ is finite, in the large $L$ limit we shall neglect $e^{-k_{a}L}$ and therefore the l.h.s. in both the equations is $e^{-k_{b}L}\approx e^{-L\sqrt{\lambda/D_{e}}}$. On the other hand, for small $\lambda$, $a^{(3)}\approx 1-\frac{v}{\omega\sqrt{D_{e}}}\sqrt{\lambda}$, whereas $a^{(1)}\approx-\frac{1+v/\sqrt{2\omega D_{e}}}{1-v/\sqrt{2\omega D_{e}}}+O(\lambda)$. This implies that the corrections in the r.h.s. of Eqs. (41)-(42) are of $O(\sqrt{\lambda})$, and the equations to the first subleading order reads,
@@
One can solve the above equation perturbatively by considering $\lambda_{n}=\lambda_{n}^{(0)}+\lambda_{n}^{(1)}$, where $\lambda_{n}^{(0)}=-\frac{n^{2}\pi^{2}D_{e}}{L^{2}}$ are the leading order solution (Eq. (45)) and $\lambda_{n}^{(1)}\ll\lambda_{n}^{(0)}$ are the corrections:
$\lambda_{n}^{(1)}\approx-\frac{2\,v^{2}}{\omega\,\sqrt{2\omega D_{e}}}\frac{\lambda_{n}^{(0)}}{L}.$ (48)
The steps are shown in Appendix C. Consequently, the band near $\lambda=0$ takes the form,
@@
We have already mentioned that $2\omega$ is an eigenvalue of the evolution operator that belongs to the symmetric sector. Let us consider $\lambda=-2\omega+\varepsilon$, where $\varepsilon$ is presumably small in the large $L$ limit. In this limit,
$k_{a}\approx\sqrt{\frac{v^{2}-2\omega D}{D^{2}}+O\left(\frac{\varepsilon}{D}\right)},~k_{b}\approx\sqrt{\varepsilon/D}$
@@
Since $\varepsilon\to 0$, $a^{(3)}\approx-1$. Therefore, for large $L$, Eqs.  (41)-(42) reduce to  [^8] Here we disregard the particular case of $v^{2}=2\omega D$.
@@ 4.3 Crossover behaviour of the relaxation time
Figure 6: Crossover behaviour of the relaxation rate for different system sizes. Here, $D=1.0,v=1.0,\omega=0.1$. Points are obtained by numerically solving Eq. (41) (symmetric sector) and identifying the solution closest to zero using Mathematica. The crossover to $L$-dependent relaxation is seen to occur for system sizes very close to $L_{c}\sim\frac{v}{\omega}=10.0$. The large $L$ behaviour is consistent with Eq. (49) for $n=1$.
For passive particles the relaxation time because of diffusion is proportional to $L^{2}/D$. RTPs on the other hand also carries a tumble dynamics which relaxes in a timescale $\tau_{t}\sim(2\omega)^{-1}$. When the system size is much larger than $L_{c}\sim v/\omega$, several tumble events occur before reaching the steady state and the relaxation dynamics becomes effectively passive-like. In this regime the bands are also well-separated. As $L$ is reduced, the corrections due to activity (Eq. (49)) become significant and a deviation from diffusive behaviour shows up. When $L\sim L_{c}$, the two bands would mix and for $L<L_{c}$ they cross, implying that the diffusive dynamics now relaxes faster compared to the tumble dynamics and $\lambda=-2\omega$ determines the relaxation of the system. This behaviour is shown in Figure 6. We expect this to be a generic feature of finite-sized active systems, and more generally of processes with coloured noise.
@@ 4.4 Eigenfunctions and the large-time distribution
Once we have the spectrum, we can completely solve for the coefficients $\beta_{n}^{i},i=(1,\dots,4)$. For $o_{x}=\pm 1$, the coefficients are related as follows:
$\beta^{(1)}$ $=$ $-\,\beta^{(3)}\,\frac{a^{(3)}\pm e^{k_{b}\,L}}{a^{(1)}\pm e^{k_{a}\,L}},$ (52)
$\beta^{(2)}$ $=$ $\pm\,\beta^{(1)}a^{(1)}\,e^{k_{a}\,L},$ (53)
$\beta^{(4)}$ $=$ $\pm\,\beta^{(3)}a^{(3)}\,e^{k_{b}\,L},$ (54)
$\beta^{(3)}$ is left arbitrary at this point. After incorporating the coefficients in Eq. (28), the expression of the eigenfunctions $|\phi_{n}(x)\rangle$ becomes,
[…]
The set of equations (52)-(54) can be simplified and the coefficients can be determined order by order in the large $L$ limit. We shall present results for the band near $\lambda=0$ and keep the terms which are $\sim$ $O(L^{-1})$. The calculation steps are given in the Appendix D. Using the coefficients in Eq. (29), we find, at large times $t\gg\,\omega^{-1}$,
$|P_{\rm tr}(x,t)\rangle$ $\approx$ $\sum_{n=1}^{\infty}c_{n}\,e^{\lambda_{n}t}\beta_{n}^{(3)}\bigg(2\,i\,\sin\frac{n\pi x}{L}\begin{bmatrix}1\\ 1\end{bmatrix}-\,i\,\frac{n\pi v}{\omega\,L}\cos\frac{n\pi x}{L}\begin{bmatrix}1\\ -1\end{bmatrix}$ (56)
@@
with $\lambda_{n}\approx-\frac{n^{2}\pi^{2}D_{e}}{L^{2}},~n>0$; here we have used $a_{n}^{(1)}\,a_{n}^{(2)}=1$. Each term of the summation in Eq. (56) is an eigenfunction evaluated to $O(L^{-1})$. In the following we shall use $\beta_{n}$ instead of $\beta_{n}^{(3)}$. To the leading order, each of the eigenfunctions is simply proportional to $\sin\frac{n\pi x}{L}$ and the probability density $P_{\rm tr}(x,t)=[1~~1]\,|P_{\rm tr}(x,t)\rangle$ takes a passive-like form with an effective diffusion constant $D_{e}$:
$\lim_{t\rightarrow\infty}\lim_{L\rightarrow\infty}P_{\rm tr}(x,t)\approx\sum_{n=1}^{\infty}c_{n}\beta_{n}\,e^{\lambda_{n}t}4\,i\,\sin\frac{n\pi x}{L}\equiv P_{\rm passive}(x,t)|_{D\rightarrow D_{e}},$ (57)
while the magnetisation vanishes. In particular, the relaxation mode is given by $\phi_{1}(x)\propto\sin\frac{\pi x}{L}$. Note that the active contributions in excess to the effective passivelike contributions, proportional to the active speed $v$, appear at $O(L^{-1})$ in each term of Eq. (56). Thus at very large times when only very few $n$’s contribute, the active part in the distribution and the magnetisation $Q_{\rm tr}(x,t)=[1~-1]\,|P_{\rm tr}(x,t)\rangle$ appears only at the subleading order. However, this is not true in the large but intermediate times when larger $n$’s also contribute.
To determine the full distribution we need to evaluate the $n$-dependent constants $\beta_{n}$ and $c_{n}$. Let $\langle\psi_{n}(x)|$ be the left eigenvenctor of the evolution operator $\mathcal{L}$ corresponding to eigenvalue $\lambda_{n}$. Then the orthonormality of the $\langle\psi_{n}(x)|$’s and $|\phi_{n}(x)\rangle$’s give us $\beta_{n}$. The $c_{n}$ is subsequently determined from the initial condition: $c_{n}=\int dx\,\langle\psi_{n}(x)|P(x,0)\rangle$. We shall consider a fixed initial condition, $|P(x,0)\rangle=\delta(x-x_{0})\,\begin{bmatrix}s_{+}\\ s_{-}\end{bmatrix}$, and find the propagator at large times. Here $s_{+},s_{-}$ are the probabilities of the initial spin to be plus and minus respectively: $s_{+}+s_{-}=1$, and $m_{0}=s_{+}-s_{-}$ is the initial magnetisation. The detailed expressions for $\langle\psi_{n}(x)|$, $\beta_{n}$ and $c_{n}$’s are given in the Appendix D. Incorporating these, we obtain the expression of the state to $O(L^{-1})$ for times $t\gg\omega^{-1}$,
@@
[…]
At relaxation time scales ($t\gtrsim L^{2}/D_{e}$), only the small $n$ terms ($n=1,2$ etc.) survive. Here the dominating contribution comes from the first term of Eq. (58) which is just the effective passive expression, while the active contribution occurs as subleading corrections. However, at intermediate times $\omega^{-1}\ll t\ll L^{2}/D_{e}\,$, all the $n$’s contribute and it is not obvious whether the active contributions constitute only a correction to the effective passive behaviour. For this we aim to evaluate the summations in a closed form. Generally this is not doable for the summations at hand, but we can make progress for large values of $L$ using the Poisson summation formula and the symmetries of the distribution. The details and a closed form expression of $|P_{\rm tr}(x,t)\rangle$ is given in Appendix D. For $x,x_{0}\ll L$ the system is well approximated by a semi-infinite line, for which we discuss the results below.
@@ 4.5 Semi-infinite line with absorbing barrier at $x=0$:
On the semi-infinite geometry the summations in the Eq. (58) can be converted to integrals and the state vector at large times takes the form:
@@
[…]
The same expression is obtained by taking the limit $L\gg x,x_{0}$ in Eq. (D.13). All the terms in the above expression are $\sim O(t^{-3/2})$, and therefore ‘passive’ as well as the ‘active’ contributions are equally important in determining the large time behaviour in the semi infinite domain. From Eq. (59) we find the distribution and magnetisation profile by respectively adding and subtracting the rows:
@@
The profiles carry an interesting structure. In each of Eqs. (60)-(61), the terms that fall exponentially in distance represents the kinetic boundary layer with a ‘skin depth’ $k_{a}^{-1}$, identical to that in the steady state. The other terms form the ‘scaling’ part of the profile.
@@
which is finite throughout the system for any finite $x_{0}$, and can actually attain dominant role for highly persistent (small $\omega$ or large $v$) run and tumble motion. The effect of activity becomes more prominent when the particle is in the boundary layer $x\lesssim k_{a}^{-1}$. Beyond that this part decays slowly as $x^{-1}$, and asymptotically the contribution from initial position $x_{0}$ only remains. In fact, it can be shown from Eq. (60) that, when both $x$ and $t$ is very large such that $y=x/\sqrt{D_{e}t}$ is finite, both the active and passive contributions are $\sim(D_{e}t)^{-1}$, and their ratio becomes,
$R(x_{0})=\lim_{\{x,t\rightarrow\infty,\frac{x}{\sqrt{D_{e}t}}=y\}}\frac{P^{\infty}_{\rm active}}{P^{\infty}_{\rm passive}}=\frac{v}{2\,\omega}\big(m_{0}+\frac{v}{\sqrt{2\,\omega\,D_{e}}}\big)\frac{1-e^{-k_{a}x_{0}}}{x_{0}}.$ (63)
Figure 8: $R(x_{0})$ from Eq. (63) as a function of $1/\omega$ for (a) $m_{0}=-1$, (b) $m_{0}=0$, (c) $m_{0}=1$ for different values of $x_{0}$. Here v=1,D=1.